Consider the process of randomly choosing 2 people from a group of 4 men and 2 women (6 people). Suppose that this process is repeated 10 times. Determine the probability that both women are chosen in at least 2 of the 10 repetitions.
Answer #1

P(both women chosen in a repetition)=2C2*4C0/6C2 =1/15

hence P(both women chosen in at least 2 of the 10 repetiions)=P(X>=2)=1-P(X<=1)=1-(P(X=0)+P(X=1))

=1-(10C0(1/15)0(14/15)10+10C1(1/15)1(14/15)9) =0.1401